OpenStudy (anonymous):

The volume (in L) that would be occupied by 5.00 mol of O2 at STP is A. 0.411 L. B. 22.4 L. C. 41.6 L. D. 112 L.

OpenStudy (anonymous):

I am not a Chemist, but I took Chemistry and passed :-D. \[P\times V=n\times R \times T\] Substitute the given values. P is pressure V is volume n is the number of moles of molecules T is the absolute temperature R is the gas constant (.0821) I maybe right or wrong about .0821 being the gas constant...

OpenStudy (anonymous):

There is also another way to solve this ^_^. 1 mole occupies 22.4 L at STP. \[5.00 \times 22.4 = 112L\] Please note: \[5.00~moles \times 22.4 L/mole = 112\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Do you understand what I was trying to say? Chemistry is sometimes confusing :-P, so I just want to check.

OpenStudy (anonymous):

yes I do